Calculate the number of hot dogs that he has to sell to break even (make a zero)
He pays the cart owner R75 per night for the use of the cart, and he sells hot dogs for R16.50 each, his cost for the hot dogs is R9. 60 per hotdog
1
Expert's answer
2020-06-22T18:13:38-0400

We can write ,

fixed cost = FC

price = P

variable cost = VC

the break even point = BE

Given in question,

FC = R75

P=R16.50

VC=9.60

BE=?

Now according to algebric formula

BE = FC/(P-VC)

BE=75/(16.50-9.60)

BE=75/6.90

BE=10.86

the number of hot dogs that he has sell to break even (make a zero ) = 10.86

0r we can say

11 hot dogs he has to sell to break even.